OpenStudy (anonymous):

y=5x+3 Use substitution to determine whether the point is on the line.

hero (hero):

What? Where's the point?

OpenStudy (anonymous):

Oops. The point is A (1,8)

OpenStudy (anonymous):

well 1st u insert the x value to see if u get the y value or not

hero (hero):

It works

hero (hero):

The point is on the line

OpenStudy (anonymous):

Ok. If it works does that mean its a line?

OpenStudy (anonymous):

y=5x+3 y=5(1)+3 y=8 therefore it duz lie on the line

hero (hero):

That means, the point is on the line.

OpenStudy (anonymous):

Ok. Thanks

hero (hero):

You might want to invest in a graphing calc

hero (hero):

so you can SEE what's going on

OpenStudy (anonymous):

ok. I will